OpenStudy (anonymous):

anyone can help me ..please..how to solve ∫(u-1)/(4u-2u²)

OpenStudy (shubhamsrg):

take 2 common out from denominator, then you'll see if you let denominator = t , then numerator = dt .. you'll have to make slight modifications just..

OpenStudy (anonymous):

Use Partial Fractions as follows \[ \frac{u-1}{4 u-2 u^2}=\frac{1-u}{2 (u-2) u}=-\frac{1}{4 u}-\frac{1}{4 (u-2)} \]

OpenStudy (anonymous):

okay... thank you @eliassaab

OpenStudy (anonymous):

yw